Consider the market for wheat where demand is given​ by:

Qd=100−1p

and supply is given​ by:

Qs=40 + 1p.

Now suppose​ that, due to a market failure​ (an artificial shipping​ constraint), a maximum of 65.00 units of wheat can be supplied by firms in the market.

The amount of the deadweight loss caused by the market failure is. ​(Enter your answer rounded to the nearest penny and as a positive

number.​)
Answer #1

ANswer

the market is in equilibrium at

Qd=Qs

100-P=40+P

2P=60

P=30

Q=100-30=70 units

======

after failure

Qs=65

price is from the demand curve at the output

65=100-P

P=35

after failure, the price is 65

but the willingness to accept of the producer is

65=P+40

P=25

=========

DWL =0.5*(the difference between willingness to accept of producer and the price after failure)*change in quantity

=0.5*(35-25)*5

=25

the DWL is $25

==========

It is also denoted by the purple color in the graph

(0, 100) 1000 Price 80+ Q=65 supply E after failure 40+ (65, 35) DWL Equilibrium before failure DWLO (70, 30) Equilibrium bef